### v1

@@ Abstract
Customization generation techniques have significantly advanced the synthesis of specific concepts across varied contexts. Multi-concept customization emerges as the challenging task within this domain. Existing approaches often rely on training a Low-Rank Adaptations (LoRA) fusion matrix of multiple LoRA to merge various concepts into a single image. However, we identify this straightforward method faces two major challenges: 1) concept confusion, which occurs when the model cannot preserve distinct individual characteristics, and 2) concept vanishing, where the model fails to generate the intended subjects. To address these issues, we introduce LoRA-Composer, a training-free framework designed for seamlessly integrating multiple LoRAs, thereby enhancing the harmony among different concepts within generated images. LoRA-Composer addresses concept vanishing through Concept Injection Constraints, enhancing concept visibility via an expanded cross-attention mechanism. To combat concept confusion, Concept Isolation Constraints are introduced, refining the self-attention computation. Furthermore, Latent Re-initialization is proposed to effectively stimulate concept-specific latent within designated regions. Our extensive testing showcases a notable enhancement in LoRA-Composer’s performance compared to standard baselines, especially when eliminating the image-based conditions like canny edge or pose estimations. Code is released at https://github.com/Young98CN/LoRA_Composer.
Keywords:
Multi-Concept Customization LoRA Integration Training-Free Controllable Generation
@@ 1 Introduction
Figure 1: Our method distinguishes itself from Mix-of-Show[7] by eliminating the image-based conditions (the sketch shown above) and the requirement to train a LoRA fusion matrix. Furthermore, we highlight the limitations of Mix-of-Show through the demonstration of failure cases. In the top row, we illustrate two key issues: concept vanishing, marked by the absence of intended concepts in the image, and concept confusion, where the model mistakenly merges and confuses distinct concepts.
@@
Diffusion models [26] have significantly advanced the field of image generation, particularly in creating images that adhere to user-specific concepts. The progress made in customization models [7, 13, 30, 27, 40, 4] play an important role in enriching the landscape of image synthesis. As technologies for single concept customization evolve, users are presented with various methods to personalize content, ranging from fine-tuning U-Net[27, 13], modifying text embeddings[5, 18], to leveraging Low-Rank Adaptations (LoRA)[10]. LoRA, in particular, serves as a versatile, plug-and-play module that enables users to customize their models to generate diverse and lifelike personal images. Its adaptability and accuracy in image generation have established LoRA as a preferred method for customization tasks, significantly influencing how the community approaches the creation of tailored visual content.
@@
While LoRA excels in single-concept customization, its application to emerging multi-concept customization tasks presents challenges. Recent developments have explored the integration of multiple-concept LoRAs to infuse images with diverse concepts via fusion tuning [7, 36]. However, as illustrated in Fig. 1, these integration strategies often necessitate a variety of conditions, including textual and image-based inputs [42] (such as human pose estimations, and canny edge detection). This approach introduces constraints on variation and flexibility. Furthermore, in the process of combining multiple LoRAs, prior research [7, 36, 32] focused on training a fusion ratio matrix. However, adjusting LoRA weights in this manner can exacerbate two problems: 1) concept vanishing, where the concept fails to be injected into the figure; and 2) concept confusion, where the model struggles to associate attributes with subjects or fails to capture distinct concept characteristics. Examples illustrating the issues mentioned are displayed in the top row of Fig. 1, showcasing outputs from the representative method Mix-of-Show[7]. The left column shows a clear case of concept vanishing, where the model fails to generate one of the anime girls and one of the dogs (highlighted within the red box). The right column highlights issues with incorrect attribute binding, such as the dog’s color being mistaken, and the anime figure’s characteristics not being distinctly expressed (indicated within the blue box).
@@
To overcome existing challenges, we introduce LoRA-Composer, a training-free framework that enables the synthesis of images with multiple integrated concepts, utilizing textual and layout cues. LoRA-Composer encompasses three principal components: Concept Injection Constraints, Concept Isolation Constraints and Latent Re-initialization. The Concept Injection Constraints introduce a novel cross-attention mechanism, consisting of: 1) Region-Aware LoRA Injection, which injects concept-specific LoRA features into designated regions through cross-attention, facilitating the seamless integration of multiple LoRAs within the diffusion model framework without the need for fusion fine-tune. 2) Concept Enhancement Constraints, which guide the refinement of latents to accentuate concepts in user-specified regions. This strategies enable the model to concentrate on areas designated for concept insertion, effectively mitigating the issue of concept vanishing. The Concept Isolation Constraints specifically address the issue of concept fusion by concentrating on self-attention. They implement restrictions to guarantee that each concept maintains its unique characteristics. This plays a crucial role in guiding the update process within the denoised latent space, maintaining alignment with each concept’s unique features and corresponding textual descriptions. Due to traditional single-concept LoRA training typically covers the entire image generation, which can be restrictive for localized area generation. We propose re-initializing the latent vector to establish a more accurate prior, directing the model’s focus on specific areas of the image.
@@
We subjected our LoRA-Composer to rigorous testing across a broad spectrum of multi-concept customization scenarios, spanning categories from pets to characters and scenic backgrounds. Our approach displayed a strong performance compared to existing benchmarks through comprehensive qualitative and quantitative assessments.
In summary, our contributions are as follows:
@@
  • We propose a training-free model for integrating multiple LoRAs called LoRA-Composer. It requires only readily accessible conditions: layout and textual prompts. This approach simplifies the process of combining various concepts into a cohesive image, enhancing the convincing for users.
@@
  • Our extensive evaluations reveal that our method exceeds baseline performance, particularly in scenario that eliminate image-based conditions.
@@ 2.1 Controllable Image Generation
Diffusion models [9, 33] trained on large-scale text-to-image datasets, like DALLE-2 [25], Imagen [29], and Stable Diffusion [26], SDXL [23] can produce text-aligned and diverse images in unprecedented high quality. To further support image generation from fine-grained spatial conditions, like sketches, human keypoints, semantic maps, etc., ControlNet [42] finetunes a trainable copy of the pre-trained U-Net and connects the new layers and original U-Net weights with zero convolutions. A similar work, T2I-Adaptor [20], finetunes lightweight adaptors for conditional generation from spatial conditions. Differently, GLIGEN [15] considers controllable generation with sparse box layout conditions and injects a gated self-attention for fine-tuning. While these works rely on self-collected datasets of condition-image pairs for finetuning, recent works [39, 22] seek to explore test-time optimization for zero-shot controllable generation. For example, both BoxDiff [39] and Attention Refocusing [22] achieve zero-shot layout conditioned generation, by maximizing the attention weights between the features inside the box and its corresponding text description, while discouraging the latent features outside the box from attending to the text.
2.2 Single-Concept Customization
Concept customization aims at generating concepts specified by a few input images. Several represented methods [27, 5, 34, 31, 37, 8, 21] focus on the customization of a single concept. For example, Dreambooth fine-tunes the pre-trained diffusion model on a few samples of the custom concept, and introduces a prior-preservation loss to alleviate overfitting. Differently, Textual Inversion [5] fixes the pre-trained weights and only tunes a learnable text embedding to bind the novel concept. To further improve the capacity of the concept representation, P+ [35] and NeTI [1] learn a layer-wise text embedding or an implicit MLP for concept binding, respectively. Another line of work [14, 28, 6, 11, 37] makes attempts to train a tailored model on large-scale image datasets for fast customization. For example, BLIP-Diffusion [14] takes both the text prompt and subject image as inputs and learns on significant amount of data pairs to generate the customized images. Differently, HyperDreamBooth [28] first collects Low-Rank Adaptation (LoRA) [10] weights of massive custom concepts, then trains the model to predict the LoRA weights from input concept image.
@@ 2.3 Multi-Concept Customization
Different from the above work that focuses on single-concept customization, several recent works [13, 16, 17, 2, 38] tackle the more challenging generic setting of multiple concepts. A pioneer work, Custom Diffusion [13], jointly finetunes multiple concept images for customization. Cones [16] finds concept-related neurons in pre-trained diffusion models for multi-concept customization. Cone 2 [17] further incorporates residual token embedding and layout conditioning for better generation quality. To accelerate the customized generation, FastComposer [38] finetunes diffusion model on massive data to take subject embedding as input and generate the composed image of multiple concepts. Similarly, Paint-by-Example [40] and AnyDoor [4] are trained on a significant amount of images and can achieve multi-concept generation through image inpainting. Considering the widespread utilization of LoRA for customization , several recent works [7, 36, 30, 44] seek to achieve multi-concept customization by combining multiple LoRA weights of individual concepts. For example, Mix-of-show [7] proposes gradient fusion to train a composed LoRA weight that mimics the prediction of individual LoRAs. It further leverages T2I-Adaptor [20] and sketches or keypoints for final generation. A concurrent work [44] proposes two variants, namely LoRA Swich and LoRA Composite, to realize LoRA merge during decoding. The former uses multiple LoRA weights sequentially, while the latter averages the latent obtained from different LoRA weights. However, they focus on the combination of a single foreground and a single background LoRA weights. By contrast, we tackle the more challenging task of customizing multiple foreground characters, facing severe issues of concept confusion and vanishing.
@@ 3 Method
Figure 2: (a) LoRA-Composer utilizes textual, layout, and image-based conditions (optional) to integrate and customize multiple concepts through Latent Re-initialization for precise layout generation. (b) Modifications to the Stable Diffusion U-Net in LoRA-Composer Block include concept isolation in self-attention and Concept Injection in cross-attention, optimizing for accurate concept placement while preventing feature leakage across concepts.
In this section, we commence with a concise overview of diffusion models and ED-LoRA in Sec. 3.1. Following this, we introduce our innovative LoRA-Composer approach in Sec. 3.2, with the pipeline depicted in Fig. 2(a). The key point is augmenting the scalability of LoRA through the utilization of the LoRA-Composer Block, as illustrated in Fig. 2(b). We then delve into the specifics of the two primary components of the LoRA-Composer Block, outlined in Sec. 3.3 and Sec. 3.4, respectively. Finally, in Sec. 3.5 we discuss the implementation of Latent Re-initialization to achieve a more refined layout generation prior. We emphasize integrating LoRAs to enable multi-concept customization within a single image, aiming for a solution that is both more flexible and scalable.
3.1 Preliminary
$\mathcal{L}=\mathbb{E}_{z\sim\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau(y))\|_{2}^{2}\right],$ (1)
where $\epsilon_{\theta}$ is the denoising unet with learnable parameter $\theta$.
ED-LoRA[7] aims to augment the expressiveness of the embedding by employing a decomposed structure. ED-LoRA implements a layer-wise embedding strategy, following the method described in P+ [35], to forge a multi-faceted representation for the concept token ($V=V_{rand}V_{class}$). This involves adding a random variation ($V_{rand}$) and a class-specific component to the base embedding ($V_{class}$). Furthermore, it integrates a LoRA layer into the linear layers present in all attention modules of the text encoder and U-Net. This integration allows for a flexible adaptation of the model to specific concepts by modifying the linear layers in a low-rank manner, thereby enhancing the model’s ability to encode and synthesize images based on textual descriptions with high fidelity. We use it by default in the following description.
@@ 3.2 LoRA-Composer Pipeline Overview
LoRA-Composer utilizes a standard LoRA approach for subject registration, facilitating seamless integration of diverse subjects without requiring retraining for LoRA fusion. Our approach to the intricate task of multi-subject customization unfolds in two steps: initially, we create an accurate representation of each subject, followed by their coherent combination into an image. The process begins by acquiring a LoRA model, which can be achieved through training or downloading one shared by the community. Following this, LoRA-Composer efficiently combines these acquired LoRAs into a unified, coherent image. Additionally, to further refine the model’s capability in managing multiple conditions simultaneously, we provide the option to incorporate image-based conditions, using T2I-Adapter[20] or ControlNet[42].
Our primary contribution is the introduction of the LoRA-Composer Block, depicted in Fig. 2(b). In this innovation, we have re-designed the attention block within the U-Net architecture. Specifically, we preserved the residual block structure while adapting both the self-attention and cross-attention mechanisms. In the self-attention layer, we introduce Concept Isolation Constraints to effectively segregate different concepts, ensuring their distinctiveness. Concurrently, within the cross-attention mechanism, we implement Concept Injection Constraints designed to counteract concept vanishing. These strategies enable the refinement of the latent space into an image customized according to user preferences, utilizing both self-attention and cross-attention maps to direct this process effectively.
@@ 3.3 Concept Injection Constraints
Simply using text prompts to specify desired concepts may result in missing concepts in the basic Stable Diffusion[3]. Although spatial attention guidance methods like BoxDiff[39], Attend-and-Excite[3], and Local control[43] can mitigate the issue of missing objects in multi-concept generation, they fall short in precisely represent user-defined concepts. To tackle the challenge of accurately incorporating user-defined concepts, we introduce the concept injection constraint approach, which is comprised of two key components: Region-Aware LoRA Injection and Concept Enhancement Constraints.
Figure 3: Modules of LoRA-Composer Block: (a) region-aware LoRA injection, (b)layout condition, (c) concept region mask.
Region-Aware LoRA Injection: Drawing inspiration from Regionally Controllable Sampling[7], our method initiates with Region-Aware LoRA Injection. As illustrated in Fig. 3(a), upon receiving a layout condition,as shown in Fig. 3(b), we extract the queries, keys, and values.
$Q_{n}=M_{n}\odot W_{0}^{Q}(z),K_{n}=W_{n}^{K}(\tau_{n}(P^{n})),V_{n}=W_{n}^{V}(\tau_{n}(P^{n})),$ (2)
@@
where $n\in[1,2,...]$ and $P^{n}$ denote the index of foreground concept LoRAs and the local prompt in a certain area respectively. Specifically, when $n=0$, $P^{n}$ indicates global prompt. The symbol $\odot$ denotes the Hadamard product, which identifies latent areas corresponding to the layout mask $M_{n}$. The symbols $W^{Q},W^{K}$, and $W^{V}$ stand for the projection matrices within the cross-attention modules of U-Net blocks which combined the concept LoRA, respectively. After replicating this process for each concept, we then update the hidden state through the cross-attention mechanism as follows:
$h_{n}=\text{softmax}\left(\frac{Q_{n}K_{n}}{\sqrt{d}}\right)V_{n},$ (3)
@@
where the $d$ represents the dimension of queries and keys. This injection approach ensures a comprehensive integration of both background and foreground concepts, enhancing the model’s ability to accurately reflect user-specified concepts within the generated images.
Concept Enhancement Constraints: Building on the methodology presented by BoxDiff[39], to ensure that synthesized objects accurately align with user-defined locations, a direct method involves confining high responses within the cross-attention mechanism to specified mask regions. Additionally, to address the issue of object generation tending towards the corners of the layout box, we incorporate a Gaussian weight into the cross-attention map sampling.
$\mathcal{L}_{ce}=\sum_{n=1}^{n}(1-\frac{1}{S}\sum^{S}_{S=0}\textbf{topk}(A^{c}_{n}\odot M_{n}\odot G,S)),$ (4)
@@
where $\textbf{topk}(\cdot,S)$ indicates the selection of $S$ elements with the highest response in input. $A^{c}_{n}$ represents the position $c$ of the concept token in the cross-attention map $A$ for a specific concept. $G$ means standard Gaussian distribution. Moreover, we also propose the loss at the projection of the x-axis and y-axis, respectively, to ensure the concepts fill the entire box. First, we squeeze cross-attention maps on the w-axis and h-axis via the max operation as below:
$a^{c}_{n}(j)=\mathbf{Max}_{i=1,2,\dots,h}{A^{c}_{n}(i,j)},$ (5)
$a^{c}_{n}(i)=\mathbf{Max}_{j=1,2,\dots,w}{A^{c}_{n}(i,j)}.$ (6)
@@
Then we compute the L1 loss in each axis as follows:
$\mathcal{L}_{fill}=\sum_{n=1}^{n}\frac{1}{K}(\sum(\mathbf{1}-\{M_{n}\odot a^{c}_{n}(j),M_{n}\odot a^{c}_{n}(i)\})),$ (7)

### v2

@@ Abstract
Customization generation techniques have significantly advanced the synthesis of specific concepts across varied contexts. Multi-concept customization emerges as the challenging task within this domain. Existing approaches often rely on training a fusion matrix of multiple Low-Rank Adaptations (LoRAs) to merge various concepts into a single image. However, we identify this straightforward method faces two major challenges: 1) concept confusion, where the model struggles to preserve distinct individual characteristics, and 2) concept vanishing, where the model fails to generate the intended subjects. To address these issues, we introduce LoRA-Composer, a training-free framework designed for seamlessly integrating multiple LoRAs, thereby enhancing the harmony among different concepts within generated images. LoRA-Composer addresses concept vanishing through concept injection constraints, enhancing concept visibility via an expanded cross-attention mechanism. To combat concept confusion, concept isolation constraints are introduced, refining the self-attention computation. Furthermore, latent re-initialization is proposed to effectively stimulate concept-specific latent within designated regions. Our extensive testing showcases a notable enhancement in LoRA-Composer’s performance compared to standard baselines, especially when eliminating the image-based conditions like canny edge or pose estimations. Code is released at https://github.com/Young98CN/LoRA_Composer
@@ 1 Introduction
Diffusion models [1] have significantly advanced the field of image generation, particularly in creating images that adhere to user-specific concepts. The progress made in customization models [2, 3, 4, 5, 6, 7] play an important role in enriching the landscape of image synthesis. As technologies for single concept customization evolve, users are presented with various methods to personalize content, ranging from fine-tuning U-Net [5, 3], modifying text embeddings [8, 9], to leveraging Low-Rank Adaptations (LoRA) [10]. LoRA is a versatile, plug-and-play module that enables users to customize their models to generate diverse and lifelike personal images. Its adaptability and accuracy in image generation have established LoRA as a preferred method for customization tasks, significantly influencing how the community approaches the creation of tailored visual content.
Figure 1: Our method distinguishes itself from Mix-of-Show [2] by eliminating the image-based conditions and the requirement to train a LoRA fusion matrix. Furthermore, we highlight the limitations of Mix-of-Show through the demonstration of failure cases. In the top row, we illustrate two key issues: concept vanishing, marked by the absence of intended concepts in the image, and concept confusion, where the model mistakenly merges and confuses distinct concepts.
@@
While LoRA excels in single-concept customization, its application to emerging multi-concept customization tasks presents challenges. Recent developments have explored the integration of multiple-concept LoRAs to infuse images with diverse concepts via fusion tuning [2, 11]. However, as illustrated in Fig. 1, these integration strategies often necessitate a variety of conditions, including textual and image-based inputs [12] (such as human pose estimations, and canny edge detection). This approach introduces constraints on variation and flexibility. Furthermore, in the process of combining multiple LoRAs, prior research [2, 11, 13] focused on training a fusion ratio matrix, which aims to optimally weigh individual LoRAs. However, adjusting LoRA weights in this manner can exacerbate two problems: 1) concept vanishing, where the concept fails to be injected into the figure; and 2) concept confusion, where the model struggles to associate attributes with subjects or fails to capture distinct concept characteristics. Examples illustrating the issues mentioned are displayed in the top row of Fig. 1, showcasing outputs from the representative method Mix-of-Show [2]. The left column shows a clear case of concept vanishing, where the model fails to generate one of the dogs (highlighted within the red box). The right column highlights issues with incorrect attribute binding, such as the dog’s color being mistaken (indicated within the blue box).
@@
To overcome existing challenges, we introduce LoRA-Composer, a training-free framework that enables the synthesis of images with multiple concept LoRAs, utilizing textual and layout cues. LoRA-Composer encompasses three principal components: concept injection constraints, concept isolation constraints, and latent re-initialization. The concept injection constraints introduce a novel cross-attention mechanism, consisting of 1) region-aware LoRA injection, which injects concept-specific LoRA features into designated regions through cross-attention, facilitating the seamless integration of multiple LoRAs within the diffusion model framework without the need for fusion fine-tune. 2) concept enhancement constraints, which guide the refinement of latents to accentuate concepts in user-specified regions. These strategies enable the model to concentrate on areas designated for concept insertion, effectively mitigating the issue of concept vanishing. The concept isolation constraints specifically address the issue of concept fusion by concentrating on self-attention. They implement restrictions to guarantee that each concept maintains its unique characteristics. Due to traditional single-concept LoRA training typically without layout conditions, which can be restrictive for localized area generation. We propose re-initializing the latent vector to establish a more accurate prior, directing the model’s focus on specific areas of the image.
@@
We rigorously test our LoRA-Composer across a broad spectrum of multi-concept customization scenarios, including categories such as animals, characters, and scenic backgrounds. Our approach displayed a strong performance compared to existing benchmarks through comprehensive qualitative and quantitative assessments. In summary, our contributions are as follows:
@@
  • We propose a training-free model for integrating multiple LoRAs called LoRA-Composer. It requires only easily accessible conditions: layout and textual prompts. This approach simplifies the process of multi-concept customization, enhancing convenience.
@@
  • Our extensive evaluations reveal that our method exceeds baseline performance, particularly in scenarios that eliminate image-based conditions.
@@ 2.1 Controllable Image Generation
Diffusion models [14, 15] trained on large-scale text-to-image datasets, like DALLE-2 [16], Imagen [17], and Stable Diffusion [1], SDXL [18] can produce text-aligned and diverse images in unprecedented high quality. To further support image generation from fine-grained spatial conditions, like sketches, human keypoints, semantic maps, etc., ControlNet [12] finetunes a trainable copy of the pre-trained U-Net and connects the new layers and original U-Net weights with zero convolutions. A similar work, T2I-Adaptor [19], finetunes lightweight adaptors for conditional generation from spatial conditions. Differently, GLIGEN [20] considers controllable generation with sparse box layout conditions and injects a gated self-attention for fine-tuning. Recent works [21, 22] seek to explore test-time optimization for zero-shot controllable generation. For example, both BoxDiff [21] and Attention Refocusing [22] achieve zero-shot layout conditioned generation, by maximizing the attention weights between the features inside the box and its corresponding text description, while discouraging the latent features outside the box from attending to the text.
@@ 2.2 Multi-Concept Customization
Concept customization aims at generating concepts specified by a few input images. While significant progress has been made in generating a single custom concept [5, 8, 23, 24, 25, 26, 27, 28, 29], the customized generation of multiple concepts remains challenging. A pioneer work, Custom Diffusion [3], jointly finetunes multiple concept images for customization. Cones series [30, 31] finds concept-related neurons in pre-trained diffusion models for multi-concept customization. To accelerate the customized generation, FastComposer [32] finetunes diffusion model on massive data to take subject embedding as input and generate the composed image of multiple concepts. Similarly, Paint-by-Example [6] and AnyDoor [7] are trained on a significant amount of images and can achieve multi-concept generation through image inpainting. Considering the widespread utilization of LoRA for customization , several recent works [2, 11, 4, 33] seek to achieve multi-concept customization by combining multiple LoRA weights of individual concepts. For example, Mix-of-show [2] proposes gradient fusion to train a composed LoRA weight that mimics the prediction of individual LoRAs. It further leverages T2I-Adaptor [19] and sketches or key points for final generation. A concurrent work [33] proposes two variants, namely LoRA Swich and LoRA Composite, to realize LoRA merge during decoding. The former uses multiple LoRA weights sequentially, while the latter averages the latent obtained from different LoRA weights. However, they focus on the combination of a single foreground and a single background LoRA weights. By contrast, we tackle the more challenging task of customizing multiple foreground characters, facing issues of concept confusion and vanishing.
@@ 3 Method
Figure 2: (a) LoRA-Composer utilizes textual, layout, and image-based conditions (optional) to integrate multiple LoRAs. (b) Modifications to the U-Net in LoRA-Composer Block include concept isolation in self-attention and concept injection in cross-attention. At timestep $t$, $z_{t}$ is first refined via $\mathcal{L}$ to ensure appearance consistency and prevent feature leakage, followed by the denoising process.
In this section, we introduce our innovative LoRA-Composer approach in section 3.1, with the pipeline depicted in Fig. 2(a). The key point is augmenting the scalability of LoRA through the utilization of the LoRA-Composer Block, as illustrated in Fig. 2(b). We then delve into the specifics of the two primary components of the LoRA-Composer Block, outlined in section 3.2 and section 3.3, respectively. Finally, in section 3.4 we discuss the implementation of latent re-initialization to achieve a more refined layout generation prior. We emphasize integrating LoRAs to enable multi-concept customization within a single image, aiming for a solution that is both more flexible and scalable.
@@ 3.1 LoRA-Composer Pipeline Overview
As shown in Fig. 2(a), LoRA-Composer utilizes a standard LoRA approach for subject registration, facilitating seamless integration of diverse subjects without requiring training for LoRA fusion. Additionally, to further refine the model’s capability in managing multiple conditions simultaneously, we provide the option to incorporate image-based conditions, using T2I-Adapter [19].
Our primary contribution is the introduction of the LoRA-Composer Block, depicted in Fig. 2(b). In this innovation, we have re-designed the attention block within the U-Net architecture. Specifically, in the cross-attention layers, we implement concept injection constraints designed to counteract concept vanishing. Concurrently, within the self-attention layers, we introduce concept isolation constraints to effectively segregate different concepts, ensuring their distinctiveness. These strategies enable the refinement of the latent space into an image customized according to user preferences, utilizing both self-attention and cross-attention maps to direct the denoise process effectively.
@@ 3.2 Concept Injection Constraints
Simply using text prompts to specify desired concepts may result in missing concepts in the basic Stable Diffusion [34]. Although spatial attention guidance methods like BoxDiff [21], Attend-and-Excite [34], and Local control [35] can mitigate the issue of missing objects in multi-concept generation, they fall short in precisely represent user-defined concepts. Additionally, Mix-of-Show [2] involves optimizing a combination of lesser LoRA weights to preserve the characteristics of the concepts within the pre-trained model. However, this can diminish LoRA’s capability to represent conceptual features, resulting in diminished concepts, as illustrated in Fig. 1. To tackle these challenges, we introduce the concept injection constraints, which is comprised of two key components: region-aware LoRA injection and concept enhancement constraints.
Region-Aware LoRA Injection. Our approach directly injects each LoRA through region-aware LoRA injection, thereby avoiding the issue of missing concepts caused by the fusion of LoRAs. As illustrated in Fig. 3(a), upon receiving a layout condition, as shown in Fig. 3(b), we extract the queries, keys, and values in the pre-defined layout $M_{i}$.
$Q_{i}=M_{i}\odot W_{0}^{Q}(z),K_{i}=W_{i}^{K}(\tau_{i}(P^{i})),V_{i}=W_{i}^{V}(\tau_{i}(P^{i})),$ (1)
@@
where $i\in\{0,1,...,N\}$. The pre-trained CLIP text encoder [36] combined with LoRA is represented by $\tau_{i}$. As depicted in Fig. 2(b), for $i\neq 0$, $i$ and $P^{i}$ correspond to the indices of foreground concept LoRAs and their associated local prompts, respectively. When $i=0$, these symbols refer to the background concept LoRA and the global prompt that describes the background. The symbol $\odot$ represents the Hadamard product. The $W^{Q},W^{K}$, and $W^{V}$ stand for the projection matrices within the cross-attention modules of U-Net blocks which combined the concept LoRA. After replicating this process for each concept, we then update the region’s hidden state through the cross-attention mechanism as follows:
$h_{i}=\text{softmax}\left(\frac{Q_{i}(K_{i})^{T}}{\sqrt{d}}\right)V_{i},$ (2)
@@
where the $d$ represents the dimension of queries and keys, this injection approach ensures a comprehensive integration of both background and foreground concepts, enhancing the model’s ability to accurately reflect user-specified concepts within the generated images.
Figure 3: Modules of LoRA-Composer Block: (a) region-aware LoRA injection, (b) layout condition, (c) concept region mask, self-attention in the gray area is not calculated.
Concept Enhancement Constraints. Region-Aware LoRA Injection does not sufficiently stimulate all inherent abilities of each LoRA, consequently resulting in the loss of detailed characteristics (see Fig. 6(e)). Previous methods [21, 34] have proposed enhancing activation within the cross-attention mechanism in specific regions to improve the ability of Stable Diffusion [1] to represent concepts effectively. However, these approaches are unsuitable for multi-concept customization tasks, as object generation tends toward the edges of the layout box (see Appendix Fig. 8(a)) or fails to fill it adequately (see Appendix Fig. 8(b)). To address these issues, we introduce a Gaussian weight into the cross-attention map, along with a $\mathcal{L}_{f}$ designed to ensure the fullness of elements within the box.
$\mathcal{L}_{c}=\sum_{i=1}^{N}(1-\frac{1}{S}\sum_{j\in E}\textbf{topk}(A^{j}_{i}\odot M_{i}\odot G,S)),$ (3)
@@
where $\textbf{topk}(\cdot,S)$ indicates the selection of $S$ elements with the highest response in input and $G$ means standard Gaussian distribution. We use the concept token by default ($V=[V_{rand},V_{class}]$, for more details, see Appendix A) during the training of concept LoRA. $E$ refers to the indexes of concept token in prompts. For the foreground concept $i$, the cross-attention map corresponding to the $j$-th token is represented by $A^{j}_{i}$. As for the $\mathcal{L}_{f}$, we squeeze cross-attention maps on the w-axis and h-axis via the max operation as below:
$a^{j}_{i}(w)=\mathbf{max}_{h=1,2,\dots,H}\{A^{j}_{i}(w,h)\},$ (4)
$a^{j}_{i}(h)=\mathbf{max}_{w=1,2,\dots,W}\{A^{j}_{i}(w,h)\},$ (5)
@@
where the variables $W$ and $H$ denote the width and height of the cross-attention map $A^{j}_{i}$. Then we compute the L1 loss in each axis as follows:
$\mathcal{L}_{f}=\frac{1}{L}\sum_{i=1}^{N}\sum_{j\in E}(\mathbf{1}-\{M_{i}\odot a^{j}_{i}(w),M_{i}\odot a^{j}_{i}(h)\}),$ (6)